Find the characteristics values and the characteristics functions of the sturn_liouville problem

d/dx[x(dy/dx]+(λ /x)y=0,

y1(1)=0, y1(e(2π))=0 where λ>0


1
Expert's answer
2021-11-08T17:39:35-0500

if "\lambda>0" :

"\lambda=k^2" with "k>0"

"x^2y''+xy'+k^2y=0"


an Euler equation with indicial equation:

"r^2+k^2=(r-ik)(r+ik)=0"

"y=c_1cos(klnx)+c_2sin(klnx)"


"y'=-c_1sin(klnx)\/x+c_2cos(klnx)\/x"

"y'(1)=c_2=0"


"y'(e^{2\pi})=-c_1sin(2k\pi)/e^{2\pi}=0"


"2k\pi=\pi n"

"k= n\/2"


"\lambda_n=(\ n/2)^2"

"y_n=cos( nlnx\/2)"